OpenStudy (anonymous):

Complex numbers z1=(3a+2) + j(3b-1) and z2=(b+1)-j(a+2-b). Write down the real and imaginary parts of z1 and z2. I understand the Real numbers from the complex numbers, but how do you obtain the fraction number of Re(z1) = 3a+2=5/7? and Re(z2) = b+1=5/7? i.e. where does the 5/7 come from?

OpenStudy (fibonaccichick666):

what do the 'j' stand for?

OpenStudy (fibonaccichick666):

are your j, the i?

OpenStudy (anonymous):

sorry yes, in engineering we use j instead of i.

OpenStudy (fibonaccichick666):

ah ok, like vectors?

OpenStudy (anonymous):

Yeah just to confuse things!

OpenStudy (fibonaccichick666):

that's so odd... yuck

OpenStudy (fibonaccichick666):

anywho, so the only info is this: \[z_1=(3a+2) + i(3b-1)\]\[z_2=(b+1)-j(a+2-b)\]

OpenStudy (fibonaccichick666):

correct?

OpenStudy (fibonaccichick666):

ignore that j

OpenStudy (fibonaccichick666):

and you are told to identify the \(Re(z_1)\) that's it?

OpenStudy (anonymous):

Yes. And for the first part of the question I had to find the real numbers of a and b; which were a=-3/7 & b=-2/7. but I cant work out the second part of the question which asks me to find the Real and Imaginary parts of z1 and z2.

OpenStudy (fibonaccichick666):

plug and chug?

OpenStudy (anonymous):

Tried but didnt work.

OpenStudy (fibonaccichick666):

ok let's check z_1 first

OpenStudy (fibonaccichick666):

so 3a+2=3(-3/7)+14/7

OpenStudy (fibonaccichick666):

-9+14=5 so 5/7

OpenStudy (fibonaccichick666):

follow?

OpenStudy (anonymous):

No where did you get 3z+2=3(-3/7)+14/7. Where did the 14/7 come from?

OpenStudy (fibonaccichick666):

14/7=2

OpenStudy (fibonaccichick666):

the re(z_1) is any part without an i

OpenStudy (anonymous):

Oh der to me! Thanks so much that clears it up. I should be ok from there for the Imaginary part. Thanks again!

OpenStudy (fibonaccichick666):

np just remember to use the LCD and you'll be fine

OpenStudy (anonymous):

LCD?

OpenStudy (fibonaccichick666):

the least common denominator for fractions ie. ⅔+½=4/6+3/6 =7/6

OpenStudy (anonymous):

Yes the simple things are throwing me! thanks heaps!

OpenStudy (fibonaccichick666):

haha they always do, for me it's 7*8 and 6*8, I can't ever remember them